OpenStudy (anonymous):

Use the given zero, r=5, to find all the zeros of the function. f(x) = x^3+x^2+9x+9,

OpenStudy (zepp):

You'll have to factor this out; \(\large x^3+x^2+9x+9\\x^2(x+1)+9(x+1)\\(x^2+9)(x+1)\)

OpenStudy (zepp):

Now, zeros are also called x-intercept, these are the values of x when y = 0.

OpenStudy (anonymous):

If one root is 5 then it must satisfy this equation.. So, factor will be : (x - 5) Divide the given equation by (x - 5) and then solve you will get the other roots too..

OpenStudy (zepp):

Set one side to 0; \(\large (x^2+9)(x+1)=0\) Then you'll have to solve for x \(\large (x^2+9)=0 \\ x^2=-9 \\ \sqrt{-9}=x=3i\) And \(x+1=0;x=-1\)